OpenStudy (anonymous):

(x^2/18y)-(7y/3z)+(5z^3/12x) how would you find LCD and solve

OpenStudy (anonymous):

you can find the LCD by simply multiplyin the denominators: (18y)(3z)(13x)

OpenStudy (anonymous):

then adjust each fraction accordingly

OpenStudy (anonymous):

the denominators are 18y, 3z and 12x lcd of the numbers is 18 lcd is 18xyz

OpenStudy (anonymous):

so how would I adjust the numerators

OpenStudy (anonymous):

multiply the 7y over 3z by six ?

OpenStudy (anonymous):

six xy

OpenStudy (anonymous):

ok first fractiom :- 18xy / 18y =x so we multiply the numerator by x so this becomes x^3

OpenStudy (anonymous):

sorry 18xyz / 18y = xz :- numerator becomes x^2 * xz = x^3z

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no it becomes x^3 z / 18xyz - its x^2 * xz to get the numerator

OpenStudy (anonymous):

for the second fraction do you multiply it by 6xy?

OpenStudy (anonymous):

in a similar way second fraction becomes (18xyz / 3z ) * 7y = 6xy * 7y = 42xy^2 so fraction becomes 42xy^2 / 18xyz

OpenStudy (anonymous):

yes multiply by 6xy

OpenStudy (anonymous):

you are getting it!

OpenStudy (anonymous):

what about the last one

OpenStudy (anonymous):

what would you multiply the numerator by?

OpenStudy (anonymous):

i must apoligise dumpling i made a mistake with the LCD - 12 wont go into 18 - the LCD should be 36xyz this makes the first fraction 2x^3z / 36xyz and the second = 84xy^2 / 36xyz

OpenStudy (anonymous):

oh ok so multiply the third by three, but after you combine all three would you need to simplify it

OpenStudy (anonymous):

third numerator is (36xyz / 12x) * 5z^3 = 3yz*5z^3 = 15yz^4 so fraction is 15yz^4 / 36xyz

OpenStudy (anonymous):

yes we now have to combine all 3 and simplify - its quite tedious algebra i'm afraid

OpenStudy (anonymous):

no need to simplify right, just put all under same denominator

OpenStudy (anonymous):

could you simplify the 84and 36

OpenStudy (anonymous):

no as you said you put it all under the same denominator - it cant be simplified any further so its [2x^3z + 84 xy^2 + 15yz^4)] / 36xyz

OpenStudy (anonymous):

- this was pretty messy but we got there in the end you cannot simplify terms when you have a sum like this as the numerator You cannot factori the numerator so cancellation is not possible.

OpenStudy (anonymous):

thanks you helped a lot